# Impact of dark states on the stationary properties of quantum particles with off-centered interactions in one dimension

### v1

@@ III.2 Competition between scales
Considering now the energy levels for $c=1.5>1$ [Fig. 2(a)], their asymptotic behavior at $\left|g\right|\gg 1$ hints at different degeneracies compared to the case where $c<1$. The absence of any triple degeneracy in the displayed energy range is attributed to the fact that no roots of low-lying Hermite polynomials exist close to $c=1.5$ and therefore dark states are not present for the low energy levels. High-level states with zeros near $c=1.5$ certainly exist and would be dark (or nearly dark) to this interaction, but we do not depict them.
@@
Aside from the two-level clustering exhibited by a few eigenstates (e.g. $n=1,2$), some low-energy levels are singly-degenerate (e.g. $n=0,3$ for $g>0$). When a singly-degenerate eigenstate is occupied, the two particles exhibit a bunching behavior inside the interval $(-c,c)$, becoming even more pronounced at strong interaction. A characteristic example is the ground bosonic state [Fig. 2(b)]. For very strong interaction ($g=100$), the two particles are almost entirely localized within the interval $(-c,c)$. However, when the particles reside in bosonic or fermionic states becoming doubly degenerate at large $g$, they display the opposite behavior. Namely, they are expelled further away from region $I\!I$ as $g$ further increases [first fermionic state in Fig. 2(c)]. This pattern is consistent with the one already encountered in the doubly-degenerate eigenstates at $c<1$ [Fig. 1(b), (c)].
@@
To understand the bunching manifested in singly-degenerate eigenstates, we first focus on the transition between single and double degeneracy. In that regard, the energy spectrum is investigated for variable displacement $c$, while keeping the interaction strength fixed at a large value, $g=10$ [Fig. 3]. When $c\ll 1$, bosonic and fermionic states with adjacent quantum numbers are doubly degenerate (exclusion regime). Their energies tend to the levels of non-interacting fermions at $c=0$ [blue dash-dotted lines on the left of Fig. 3]. For slightly larger $c$, the doubly degenerate energy levels increase since the particles are pushed away from the interaction centers, to the edges of the harmonic trap where the potential energy is higher [Fig. 1 (b), (c)].
@@
As $c$ further increases, there are particular points where the energy levels of bosonic states with quantum numbers $n$ and $n+2$ approach the energy of a fermionic eigenstate with number $n+1$ [e.g. dashed circle in Fig. 3]. These points mark a triple degeneracy, which will be clearly manifested as $g\to\infty$ [see Fig. 1 (a)]. The corresponding $c$ values lie very close to the first root of even Hermite polynomials, thus further establishing the link between triple degeneracy and dark states. When the displacement is tuned a bit further away from these points, the energy levels of the adjacent bosonic states anticross, and one of them becomes singly-degenerate [e.g. $n=0$ at $c\gtrsim 1$ or $n=2$ at $c\lesssim 1$]. This implies that dark states mark the onset of singly-degenerate eigenstates. Moreover, they set the boundaries between single and double degeneracy.
@@
Apart from bosonic eigenstates, the energy levels of fermionic states can also be singly-degenerate. The transition from double to single degeneracy takes place close to the roots of odd Hermite polynomials [e.g. dash-dotted circle in Fig. 3]. At these values, two adjacent fermionic levels with quantum numbers $n$ and $n+2$ become degenerate with a bosonic state corresponding to $n+1$.
@@
Close to the triple degeneracy points, doubly- and singly-degenerate energy levels coexist, in what we call the crossover regime. As the displacement $c$ however is further tuned to higher values, singly-degenerate eigenstates outnumber any other kind at low energies. These levels tend towards the non-interacting bosonic and fermionic values as $c\gg 1$ [blue dash-dotted lines on the right of Fig. 3]. In this region, the interaction centers are located at the edges of the harmonic trap, and the two particles barely feel any interaction. The oscillator length is the only relevant length scale, dictating the exponential decay of the eigenstates at large separations. Since the energy spectrum resembles that of an harmonic oscillator, truncated at the edges due to the interaction centers, the regime of large interaction displacement is called the truncation region.
@@
Close to the triple degeneracy mark however [dashed circle in Fig. 4], small variations in the displacement $c$ result in a substantial drop of the spatial extent of the ground bosonic state, almost an order of magnitude. The particles are now found at very small separations, a manifestation of the bunching effect for singly-degenerate eigenstates [see also Fig. 2 (b)]. As $c$ is tuned to further larger values, the extent of the ground state asymptotes to the non-interacting width, $\braket{x_{0}^{2}}=1$ Sakurai (1967) [horizontal blue dash-dotted line in Fig. 4]. Note that $\braket{x_{0}^{2}}$ approaches unity from below, indicating that the bunching is strong, and the particles localize within smaller distances than the oscillator length. The same abrupt drop occurs in the spatial extent of the ground fermionic state as well, in the vicinity of the first root of the $n=3$ Hermite polynomial [dash-dotted circle in Fig. 4]. From this point on, this fermionic state becomes singly-degenerate, and subsequently $\braket{x_{1}^{2}}$ saturates to the non-interacting value, 3 [black dash-dotted horizontal line in Fig. 4].
The spatial extent of higher excited states features a very interesting pattern. For the first excited bosonic state for example, $\braket{x_{2}^{2}}$ goes through a series of sudden drops and increases [$n=2$ in Fig. 4], as $c$ tends to larger values. These abrupt changes occur nearby triple degeneracy points [see also Fig. 3]. The two drops [$c\simeq 0.75,2.3$] are attributed to the single degeneracy and hence to the bunching effect. The one increase in between [$c\simeq 1$] takes place because from this point on the $n=2$ eigenstate becomes doubly degenerate with the $n=1$ fermionic state. Due to these series of abrupt transitions, the spatial extent $\braket{x_{2}^{2}}$ assumes its asymptotic value 5 [yellow dash-dotted horizontal line] at a larger displacement in comparison to the other two eigenstates. Let us also note that for even stronger repulsions, the changes in $\braket{x_{n}^{2}}$ near the triple degeneracy points grow increasingly abrupt.
IV Infinite interactions
Despite the fact that the onset of single, double and triple degeneracy can be clearly observed already from large interactions [Figs. 1, 2 and 3], the degeneracies occur only at infinite interaction strengths. Even though this is a special regime requiring separate treatment, further insight can be gained regarding the structure of eigenstates with different degree of degeneracy.
IV.1 Energy levels
@@
To find the energy level structure at $g=+\infty$, we proceed as in Sec. III. Namely, the relative $x$ coordinate is separated into three regions, $I$, $I\!I$ and $I\!I\!I$. The boundary conditions that the wave functions and their derivatives satisfy at the interval intersections are the same as for finite $g$ [Eqs. (13a)-(13d)]. The most important information that we get from Eqs. (13c), (13d) is that the wavefunctions have to vanish at the interaction centers, $x=\pm c$. This guarantees that the first derivatives remain finite.
This leads us to update the setup into three disjoint regions, given that there is effectively a hard wall at the interaction centers. The relative Hamiltonian in regions $I$ and $I\!I\!I$ corresponds to the Hamiltonian of a single particle confined in a harmonic oscillator, truncated by a hard-wall boundary at $x=-c$ ($I$) or $x=c$ ($I\!I\!I$). In region $I\!I$, $H_{\rm{rel}}$ corresponds to that of a single particle confined in a box potential superimposed with a harmonic oscillator.
We first focus on the latter region. The relative wave functions can be written as a superposition of parabolic cylinder functions [see also Eq. (11)],
where $Q^{(in)}_{n}\equiv\epsilon^{(in)}_{n}-1/2$. The superscript denotes that the two particles are located inside the interval determined by the interaction centers, $(-c,c)$. The hard-wall boundary conditions imposed by the box potential imply that $\phi^{(in)}_{n}(-c)=\phi^{(in)}_{n}(c)=0$. These two equations can be cast in the form of a matrix eigenvalue problem with zero eigenvalue Consortini and Frieden (1976),
$D^{2}_{Q^{(in)}_{n}}(c)-D^{2}_{Q^{(in)}_{n}}(-c)=0.$ (19)
For the other two symmetric regions it suffices to solve the relative Hamiltonian in only one of them. This is due to the fact that $H_{\rm{rel}}$ is invariant under spatial inversion ($\Pi$), even for infinite interactions [see also Sec. II]. As a result, the energy levels in regions $I$ and $I\!I\!I$ coincide, denoted as $\epsilon^{(out)}_{\nu}$. Focusing on region $I\!I\!I$ for example, the wave functions read $\chi^{(out,I\!I\!I)}_{\nu}(x)=\alpha^{(out,I\!I\!I)}D_{Q^{(out)}_{\nu}}(x)$, where $Q^{(out)}_{\nu}\equiv\epsilon^{(out)}_{\nu}-1/2$. The superscript denotes that both particles are located outside of the interaction centers interval $(-c,c)$, and in particular in region $I\!I\!I$. Imposing the hard-wall boundary condition at $x=c$, the energy levels $\epsilon^{(out)}_{\nu}$ are determined Grosche (1993),
$D_{Q^{(out)}_{\nu}}(c)=0.$ (20)
The associated eigenstates $\chi^{(out,I\!I\!I)}_{\nu}(x)$ however do not transform properly under the $\Pi$ and $\Sigma$ operations, since they are mapped to interval $I$, e.g. $\hat{\Sigma}\chi^{(out,I\!I\!I)}_{\nu}(x)=\chi^{(out,I)}_{\nu}(x)$. To resolve this issue adequate superpositions of the eigenstates in the two regions are constructed,
Figure 5: (a) Energy levels at $g=+\infty$ versus the displacement $c$. The positions of the exact crossings between the doubly-degenerate $\epsilon^{(out)}_{n}$ and singly-degenerate $\epsilon^{(in)}_{n}$ levels designate three regions, the exclusion (E), crossover (C), and truncation (T) region.
From the above analysis it becomes clear that the bosonic and fermionic levels $\epsilon^{(out)}_{n}$ are doubly-degenerate, and that particles occupying these eigenstates are localized outside of the interval $(-c,c)$. On the other hand, the two particles are strictly found within $(-c,c)$ when the $\phi^{(in)}_{n}$ eigenstates are populated. In this case the bosonic and fermionic levels are distinct [Eq. (19)], they are singly-degenerate. Varying the displacement $c$ is equivalent to moving the hard walls, and thus shifting the energy levels $\epsilon^{(in)}_{n}$ and $\epsilon^{(out)}_{n}$ [Fig. 5]. What distinguishes the spectra from those with large finite $g$ [Fig. 3] is that the hard walls at $x=\pm c$ prevent any transition between singly- and doubly-degenerate eigenstates. This is captured in Fig. 5, where there are only exact crossings between $\epsilon^{(in)}_{n}$ and $\epsilon^{(out)}_{n}$ for variable displacement $c$. The positions of these crossings correspond to roots of Hermite polynomials, marking the presence of dark states, and hence triple degeneracy.
At large finite interactions, the exact crossings become avoided [Fig. 3]. In this case, the double-to-single degeneracy transition manifests, identified in Figs. 3 and 4. In this regime, there are no hard walls at $x=\pm c$ [or additional symmetry, see the follow-up Section] preventing the particles from changing their spatial configuration. When $c$ lies close to the roots of Hermite polynomials, a transition occurs between a spatial structure resembling the one for $\phi^{(in)}_{n}$ (single-degeneracy) and the one similar to $\phi^{(out)}_{n}$ (double-degeneracy). From this vantage point, the bunching occurs because the density profiles of singly-degenerate eigenstates at $g\gg 1$ resemble the structure of $\phi^{(in)}_{n}$ at $g=+\infty$.
Apart from understanding the bunching effect, the energy levels at $g=+\infty$ [Fig. 5] help us also to neatly classify the density profiles of the two-particle system, depending on the value of $c$. In particular, $\epsilon^{(in)}_{0}$ and $\epsilon^{(out)}_{0,1}$ delineate the three regimes that were identified before [see also Sec. III]. On the left of $\epsilon^{(in)}_{0}$, there is the exclusion regime (E). All eigenstates are doubly-degenerate, and particles are excluded from the interval $(-c,c)$. On the right of $\epsilon^{(out)}_{0,1}$, all eigenstates are singly-degenerate, marking the truncation regime (T). Particles are localized within the interaction center interval $(-c,c)$, and the energy levels saturate to their non-interacting values at large $c$. In-between these two regions, doubly- and singly-degenerate eigenstates coexist, denoting the crossover regime (C). In this regime dark states occur at the positions of the exact crossings, a sign of triple degeneracy. Of course, such a classification scheme serves only as a guide when tackling large finite interactions. The boundaries between the regions slightly change due to the occurrence of avoided crossings [Fig. 3].
IV.2 Symmetry analysis
For finite $g$ and $c$, all energy levels of the relative sub-Hamiltonian are singly-degenerate, as is true for any solutions to the 1D Schrödinger equation defined on a path-connected interval. More generally, any system whose maximal kinematic symmetry is Abelian should only be singly-degenerate. How then can we understand the additional double and triple degeneracies that occur in the $g\to\infty$ limit?
These can be understood by recognizing that the domain of the relative coordinate $x$ becomes effectively disconnected for infinite $g$. Each of the three regions $I$, $I\!I$ and $I\!I\!I$ acts like an independent quantum system, each experiencing independent time evolution. Energy eigenstates are localized to each of these three independent regions in the $g\to\infty$ limit. To see this, note that the relative sub-Hamiltonian $H_{\mathrm{rel}}(g,c)$ can be re-expressed as the direct sum of three Hamiltonians:
$\lim_{g\to\infty}H_{\mathrm{rel}}(g,c)=H_{I}\oplus H_{I\!I}\oplus H_{I\!I\!I}$ (23)
where $H_{R}$ is the restriction of the 1D harmonic oscillator Hamiltonian restricted to region $R\in\{I,I\!I,I\!I\!I\}$. These restricted Hamiltonians commute, so instead of one time translation symmetry parameterized by $t\in T_{t}\sim\mathbb{R}$, there are now three parameterized by $(t_{I},t_{I\!I},t_{I\!I\!I})\in T_{t}^{\times 3}\sim\mathbb{R}^{3}$ and represented by products of the three unitary operators $\hat{U}_{R}(t_{R})=\exp(-iH_{R}t_{R})$. Equivalently, the phase difference between disjoint regions in the $g\to\infty$ limit is not an observable quantity Harshman (2017a). Only when $g$ is finite, is there coupling between adjacent regions that locks their relative phase.
However, this additional symmetry of tripled time evolution is still an Abelian kinematic symmetry, and therefore not enough to explain the systematic double and triple degeneracies that occur for the $g\to\infty$ limit of $H_{\mathrm{rel}}(g,c)$. If regions $I$, $I\!I$, and $I\!I\!I$ were all intervals with inequivalent domains, then the spectrum would be singly-degenerate except for so-called accidental degeneracies where two states (or even more rarely, three states) would coincide in energy. For our system, the regions $I$ and $I\!I\!I$ are equivalent intervals exchanged by relative parity $\Sigma$. Therefore the spectrum of $H_{I}$ and $H_{I\!I\!I}$ coincide and the out states $\chi_{\nu}^{(out,I)}$ and $\chi_{\nu}^{(out,I\!I\!I)}$ are double-degenerate pairs that can be symmetrized into bosonic or antisymmetrized into fermionic states $\phi_{n}^{(out)}$.
$\hat{\Sigma}\hat{U}_{I}(t_{I})=\hat{U}_{I\!I\!I}(t_{I\!I\!I})\hat{\Sigma}.$ (24)
Further, triple degeneracies occur when states in the spectrum of region $I\!I$ have energies that coincide with states of the outer regions $I$ and $I\!I\!I$. These spectral ‘accidents’ occur precisely when the displacements ±$\pm c$ align with the $n$th Hermite polynomial root, $H_{n}(c/\sqrt{2})=0$. When $n$ is even, the state in region $I\!I$ is even, and the triply-degenerate state is comprised of two bosonic states (one inside and two outside the central region) and one fermionic state. When $n$ is odd, triple degeneracy occurs between two fermionic states and one bosonic. Understanding these accidental degeneracies as the result of dark states provides an interesting perspective on an old subject Jauch and Hill (1940); McIntosh (1959); Louck and Metropolis (1981); Moshinsky and Quesne (1983); Leyvraz et al. (1997), but not all accidental degeneracies can be understood this way.
@@ V Summary and Conclusions
We have investigated the stationary properties of two harmonically trapped particles interacting via contact potentials with a displacement $c$. Depending on the value of the latter, the energy spectra are classified into three regimes. The exclusion regime for $c$ smaller than the oscillator length, where the energy levels of adjacent bosonic and fermionic states cluster together as $g\to\infty$. The two particles are found at large separations, and they are expelled from the interval dictated by the interaction centers, $(-c,c)$. In the truncation regime occurring at $c$ larger than the oscillator length, all energy levels are singly-degenerate as $c\to\infty$. The corresponding relative wavefunctions have a non-zero support only at short interparticle distances, signalling a bunching effect. This is understood in terms of the stationary properties at infinite interactions, where the relative Hamiltonian in the interval $(-c,c)$ is equivalent to that of a box potential superimposed with a harmonic oscillator. In the crossover regime, as the name suggests, both singly- and doubly-degenerate eigenstates coexist.

### v3

@@ IV.2 Competition between scales
We now consider the energy levels for $c=1.5>1$, spanning both the crossover and truncation region [Fig. 3(a)]. The absence of any triple degeneracy in the displayed energy range is attributed to the fact that no roots of low-lying Hermite polynomials exist close to $c=1.5$ and therefore dark states are not present for the low energy levels. High-level states with zeros near $c=1.5$ certainly exist and would be dark (or nearly dark) to this interaction, but we do not depict them.
@@
Aside from the two-level clustering exhibited by a few eigenstates (e.g. $n=1,2$), some low-energy levels are non-degenerate (e.g. $n=0,3$ for $g>0$). When a non-degenerate eigenstate is occupied, the two particles exhibit a bunching behavior inside the interval $(-c,c)$, becoming even more pronounced at strong interaction. A characteristic example is the ground bosonic state [Fig. 3(b)]. For very strong interaction ($g=100$), the two particles are almost entirely localized within the interval $(-c,c)$. At this regime, the non-degenerate eigenstates are similar to $\phi^{(in)}_{n}(x)$, wich are exclusively supported within the interaction center interval. However, when the particles reside in bosonic or fermionic states becoming doubly degenerate at large $g$, they display the opposite behavior. Namely, they are expelled further away from region $I\!I$ as $g$ further increases [first fermionic state in Fig. 3(c)]. This pattern is consistent with the one already encountered in the doubly-degenerate eigenstates at $c<1$ [Fig. 2(b), (c)].
@@
To further understand the transition between exclusion, crossover, and truncation regime, the energy spectrum is investigated for variable displacement $c$, while keeping the interaction strength fixed at a large value, $g=10$ [Fig. 4]. When $c\ll 1$, bosonic and fermionic states with adjacent quantum numbers are doubly degenerate (exclusion regime). Their energies tend to the levels of non-interacting fermions at $c=0$ [blue dash-dotted lines on the left of Fig. 4]. For slightly larger $c$, the doubly degenerate energy levels increase since the particles are pushed away from the interaction centers, to the edges of the harmonic trap [Fig. 2 (b), (c)] where the potential energy is higher.
@@
As $c$ further increases, there are particular points where the energy levels of bosonic states with quantum numbers $n$ and $n+2$ approach the energy of a fermionic eigenstate with number $n+1$ [e.g. dashed circle in Fig. 4]. These points mark a triple degeneracy, which will be clearly manifested as $g\to\infty$ [see Fig. 1 ]. The corresponding $c$ values lie very close to the first root of even Hermite polynomials, thus further establishing the link between triple degeneracy and dark states. When the displacement is tuned a bit further away from these points, the energy levels of the adjacent bosonic states anticross, and one of them becomes non-degenerate [e.g. $n=0$ at $c\gtrsim 1$ or $n=2$ at $c\lesssim 1$]. This implies that dark states mark the onset of non-degenerate eigenstates. Moreover, they set the boundaries between single and double degeneracy.
@@
Apart from bosonic eigenstates, the energy levels of fermionic states can also be non-degenerate. The transition from double to single degeneracy takes place close to the roots of odd Hermite polynomials [e.g. dash-dotted circle in Fig. 4]. At these values, two adjacent fermionic levels with quantum numbers $n$ and $n+2$ become degenerate with a bosonic state corresponding to $n+1$.
@@
Close to the triple degeneracy points, doubly- and non-degenerate energy levels coexist, in what we call the crossover regime. As the displacement $c$ however is further tuned to higher values, non-degenerate eigenstates outnumber any other kind at low energies. Similarly to the $g=\infty$ scenario [Fig. 1], these levels tend towards the non-interacting bosonic and fermionic values as $c\gg 1$ [blue dash-dotted lines on the right of Fig. 4].
@@
Close to the triple degeneracy mark however [dashed circle in Fig. 5], small variations in the displacement $c$ result in a substantial drop of the spatial extent of the ground bosonic state, almost an order of magnitude. The particles are now found at very small separations, a manifestation of the bunching effect for non-degenerate eigenstates [see also Fig. 3 (b)]. As $c$ is tuned to further larger values, the extent of the ground state asymptotes to the non-interacting width, $\braket{x_{0}^{2}}=1$ Sakurai (1967) [horizontal blue dash-dotted line in Fig. 5]. Note that $\braket{x_{0}^{2}}$ approaches unity from below, indicating that the bunching is strong, and the particles localize within smaller distances than the oscillator length. The same abrupt drop occurs in the spatial extent of the ground fermionic state as well, in the vicinity of the first root of the $n=3$ Hermite polynomial [dash-dotted circle in Fig. 5]. From this point on, this fermionic state becomes non-degenerate, and subsequently $\braket{x_{1}^{2}}$ saturates to the non-interacting value, 3 [black dash-dotted horizontal line in Fig. 5].
@@
The spatial extent of higher excited states features a very interesting pattern. For the first excited bosonic state for example, $\braket{x_{2}^{2}}$ goes through a series of sudden drops and increases [$n=2$ in Fig. 5], as $c$ tends to larger values. These abrupt changes occur nearby triple degeneracy points [see also Fig. 4]. The two drops [$c\simeq 0.75,2.3$] are associated to the absence of any degeneracy and hence to the bunching effect. The one increase in between [$c\simeq 1$] takes place because from this point on the $n=2$ eigenstate becomes doubly degenerate with the $n=1$ fermionic state. Due to these series of abrupt transitions, the spatial extent $\braket{x_{2}^{2}}$ assumes its asymptotic value 5 [yellow dash-dotted horizontal line] at a larger displacement in comparison to the other two eigenstates. Let us also note that for even stronger repulsions, the changes in $\braket{x_{n}^{2}}$ near the triple degeneracy points grow increasingly abrupt.
@@ V Summary and Conclusions
We have investigated the stationary properties of two harmonically trapped particles interacting via contact potentials with a displacement $c$. Depending on the value of the latter, the energy spectra are classified into three regimes. The exclusion regime for $c$ smaller than the oscillator length, where the energy levels of adjacent bosonic and fermionic states cluster together as $g\to\infty$. The two particles are found at large separations, and they are expelled from the interval dictated by the interaction centers, $(-c,c)$. In the truncation regime occurring at $c$ larger than the oscillator length, all energy levels are non-degenerate as $c\to\infty$. The corresponding relative wavefunctions have a non-zero support only at short interparticle distances, signalling a bunching effect. This is understood in terms of the stationary properties at infinite interactions, where the relative Hamiltonian in the interval $(-c,c)$ is equivalent to that of a box potential superimposed with a harmonic oscillator. In the crossover regime, as the name suggests, both singly- and doubly-degenerate eigenstates coexist.
@@
However, the primary motivation for considering this two-body model is that it is analytically tractable, one which exemplifies the role of an additional length scale mimicking a finite interaction range. Such a length scale leads to rich patterns in the energy spectrum and eigenstates, which can be classified by means of the dark states present in the system.